OpenStudy (anonymous):

How would I find. The solution of the system? Use elimination. 3x-4y=9-3x+2y=9

OpenStudy (texaschic101):

3x - 4y = 9 -3x + 2y = 9 are these your 2 equations ?

OpenStudy (anonymous):

Yes

OpenStudy (texaschic101):

then add them and what do you get ?

OpenStudy (texaschic101):

keep in mind, a -3 and a + 3 cancel....so your x's will be eliminated

DivineSolar (divinesolar):

After the png, Substitute (4y+9) for x in −3x+2y=9:

OpenStudy (texaschic101):

the question specifies elimination, not substitution

DivineSolar (divinesolar):

Yes but it could potentially be solved either way.

OpenStudy (texaschic101):

true ^^

DivineSolar (divinesolar):

Substitution would be harder but it would get the job done.

DivineSolar (divinesolar):

And after you do the steps after that, Substitute ( −18 5 ) for y in x=4y+9:

DivineSolar (divinesolar):

Those 2 will be your answers.

OpenStudy (anonymous):

A. (3,9) b. (-27, -9) c (-3,-6) d. (-9,-9) these are answers

OpenStudy (texaschic101):

3x - 4y = 9 -3x + 2y = 9 -------------add -2y = 18 y = 18/-2 y = -9 now sub -9 in for y in either of your original equations to find x 3x - 4y = 9 3x - 4(-9) = 9 3x + 36 = 9 3x = 9 - 36 3x = - 27 x = -27/3 x = - 9 you can then check your answers by subbing in what you got for x and y -3x + 2y = 9 -3(-9) + 2(-9) = 9 27 - 18 = 9 9 = 9 (correct) so your solution is : (-9,-9)